# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@
Adversarial attacks, which manipulate input data to undermine model availability and integrity, pose significant security threats during machine learning inference. With the advent of Large Vision-Language Models (LVLMs), new attack vectors, such as cognitive bias, prompt injection, and jailbreak techniques, have emerged. Understanding these attacks is crucial for developing more robust systems and demystifying the inner workings of neural networks. However, existing reviews often focus on attack classifications and lack comprehensive, in-depth analysis. The research community currently needs: 1) unified insights into adversariality, transferability, and generalization; 2) detailed evaluations of existing methods; 3) motivation-driven attack categorizations; and 4) an integrated perspective on both traditional and LVLM attacks. This article addresses these gaps by offering a thorough summary of traditional and LVLM adversarial attacks, emphasizing their connections and distinctions, and providing actionable insights for future research.
@@ 1. Introduction
[Role of AEs (section 2.1.2), draw=harvestgold, fill=harvestgold!15, text width=3.2cm, text=black]
@@
[Speed (section 5.4), draw=majorblue, fill=majorblue!15, text width=3cm, text=black]
ATKs beyond Classification (section 6), draw=brightlavender, fill=brightlavender!15, text width=3.5cm, text=black
@@
Adversarial attacks meticulously manipulate inputs to maliciously compromise model availability and integrity, posing significant security threats during machine learning inference. These attacks affect critical applications such as facial recognition (Sharif et al., 2016; Komkov and Petiushko, 2021; Zheng et al., 2023), pedestrian detection (Thys et al., 2019), autonomous driving (Eykholt et al., 2018; Chen et al., 2019; Song et al., 2018), and automated checkout systems (Liu et al., 2020), with severe implications for system security. To improve robustness and safeguard these applications, researchers have pursued extensive investigations, as demonstrated by competitions like NIPS 2017 (NIPS, 2017) and GeekPwn CAAD 2018 (GeekPwn, 2018). A comprehensive understanding of the evolution of adversarial attacks is essential for developing more effective defenses, especially in the large language models (LLMs) context. However, classical reviews often fail to capture the latest advancements (Yuan et al., 2019; Akhtar and Mian, 2018; Akhtar et al., 2021), while recent surveys tend to focus on specific areas (Wei et al., 2024b; Badjie et al., 2024; Wang et al., 2024d; Jin et al., 2024; Liu et al., 2024d) or lack thorough summaries (Costa et al., 2024). This paper differentiates itself from existing reviews in several key aspects:
Key Concepts Extraction (section 2). Adversariality, transferability, and generalization are critical traits of AEs that inform design objectives and motivations. This paper fills gaps in previous works by summarizing the causes of adversariality and transferability (section 2.1.1 and section 2.2.1), the roles of AEs (section 2.1.2), the properties of transferability (section 2.2.2), and the different types of generalization (section 2.3), which are often overlooked in existing literature.
Motivation Emphasis in Classification (section 4 and section 5). Motivation drives the achievement of goals, which often vary depending on the attacker’s knowledge level and context. As illustrated in Fig. 3, we first categorize attack methods in stage 2 based on knowledge levels and then summarize the design motivations within each knowledge context. Unlike previous works that primarily classify attacks by knowledge levels, we provide a deeper analysis of the motivations behind them.
@@
This paper provides a comprehensive overview of adversarial attack developments, with key contributions (see Fig. 1):
@@
Identifying emerging attack paradigms and potential vulnerabilities in LVLMs (section 7.1.2).
@@ 2.1.1. Why Do Adversarial Examples Exist?
Understanding the underlying reasons for AEs is essential for designing stronger attacks and more robust systems. Here, we summarize the reasons outlined in previous literature:
Linear Nature of Neural Networks (Goodfellow et al., 2014; Carlini and Wagner, 2017; Papernot et al., 2016a; Tramèr et al., 2017b; Gao et al., 2020b, a). Despite employing nonlinear activations, the linear traits of DNNs in high-dimensional space highly contribute to the existence of AEs. Consider a simple linear model defined as . Applying a perturbation , which is comparable to the model parameters , to the input can markedly alter the predictions (Goodfellow et al., 2014). This finding provided the basis for the famed FGSM (Goodfellow et al., 2014). Additionally, CW (Carlini and Wagner, 2017) demonstrated a strong correlation between interpolated samples and network logits by interpolating between the origins and AEs, thus practically supporting the linearity hypothesis.
Large Gradient Around Decision Boundaries (Papernot et al., 2016b). This indicates that small perturbations of data points can lead to significant changes in predictions, with points near decision boundaries being potential AEs. This concept also supports the motivations behind the CWA (Chen et al., 2023c) and the RAP (Qin et al., 2022), which enhance transferability by encouraging AEs to converge toward flatter regions—flatter regions contribute to better generalization (Hao et al., 2019).
@@
Additionally, (Szegedy, 2013; Sharif et al., 2016; Tabacof and Valle, 2016) suggest that AEs manifest as low-probability, high-density “pockets” within high-dimensional manifolds, rendering them difficult to acquire through random sampling. However, counterarguments exist. Some researchers (Liu et al., 2016; Dong et al., 2018) contend that the linear hypothesis may not apply when perturbations are substantial. Others have discovered that networks can also be susceptible to low-frequency information (Guo et al., 2018; Sharma et al., 2019).
2.1.2. Role of Adversarial Examples
In addition to attacking models, AEs have various other uses:
Copyright Protection (Shan et al., 2023; Dong et al., 2023b). In tasks related to infringement, such as style transfer (Johnson et al., 2016) and face swapping (Bode et al., 2021), adversarial attacks can serve to protect artists’ copyrights (Shan et al., 2023) and reduce the quality of swaps (Dong et al., 2023b).
AEs are a double-edged sword: they disrupt systems while concurrently enhancing insights and driving improvements.
@@ 2.2.2. Characteristics of Transferability
Through analysis and experiments, researchers have gained valuable insights into transferability. For example, (Kurakin et al., 2018, 2016) argue that the transferability of iterative methods is inferior to that of single-step methods, while (Zhou et al., 2018) contend that properly guided gradients can enable iterative methods to achieve good transferability. Additionally, (Kurakin et al., 2016) found that transferability may be inversely proportional to adversariality. By comparing model-aggregated samples (Kurakin et al., 2016), it was shown that universal samples can further enhance transferability (Moosavi-Dezfooli et al., 2017). Moreover, (Kurakin et al., 2016) discovered that adversarial training with highly transferable samples improves model robustness, whereas (Tramèr et al., 2017b) suggested that a higher dimensionality of adversarial subspaces leads to a greater intersection between the adversarial subspaces of two models, resulting in improved transferability.
@@ 2.3. Generalization
Cross-Environment (Physical Robustness). This generalization allows AEs to maintain their adversariality across different device environments, such as those encountered with smartphones, cameras, or printers (Kurakin et al., 2018; Eykholt et al., 2018; Thys et al., 2019). This phenomenon is often referred to as physical robustness.
@@
Cross-Corpus. This allows samples to exhibit general adversarial semantics, resulting in query-agnostic effects.
Both Cross-Prompt (Bailey et al., 2023; Lu et al., 2024; Luo et al., 2024a) and Cross-Corpus (Qi et al., 2024; Wang et al., 2024e; Ying et al., 2024b) can achieve transfer effects across prompts, but their implementation methods differ. Cross-Corpus aligns the outputs of the LVLM with malicious corpora, enabling perturbations to develop general adversarial semantics and ultimately achieve cross-prompt adversariality. In contrast, Cross-Prompt aggregates perturbations directly across a set of prompts to facilitate generalization.
@@ 3. Problem Setting
In this section, we define adversarial attacks and the threat model in section 3.1 and section 3.2, respectively. This is followed by a discussion of relevant evaluation frameworks in section 3.3, section 3.4, and section 3.5. Since some existing victim models are protected by defense strategies, section 3.3 will also address these defense strategies against adversarial attacks.
3.1. Problem Definition
Here, represents the image input modified by , and denotes the desired model output from the attacker. The goal is to identify an attack function for effective input modification. Additionally, data types such as videos and point clouds can also serve as visual inputs. The form of varies depending on the task: for classification, it yields labels; for detection, it produces bounding boxes; and for multimodal tasks, it may generate textual responses. Based on the attack intent, targets can be categorized into targeted and untargeted attacks. In targeted attacks, is constrained to be as close as possible to the attacker-specified target output ; in untargeted attacks, must be as far as possible from the original model output . The optimization paradigm for the attack can be expressed as follows:
Here, represents the modification to the input image, and denotes the distance function. For image adversarial attacks, there are two methods to constrain to ensure that the perturbations remain imperceptible: 1) Using box constraints 111The box constraint has two meanings: 1) limiting perturbed pixel values to a valid color range (Carlini and Wagner, 2017), and 2) restricting the perturbation magnitude itself. In this paper, the box constraint mainly refers to the stronger latter. to limit the size of perturbations in the pixel domain, typically implemented by constraining (where is a hyperparameter representing the budget); 2) Not constraining the size of the perturbations but ensuring that they are visually imperceptible (see UAEs in section 5.3). While this may allow for larger perturbations, it tends to create specific shapes that appear natural and undetectable to the human eye (Chen et al., 2023a).
@@ 3.2.2. Attacker Goals
Traditional taxonomies divide attack goals into targeted and untargeted attacks. Targeted attacks aim to force the model to produce predictions specified by the attacker, while untargeted attacks require only that the model’s predictions deviate from the correct output. Since targeted and untargeted attacks can be easily interchanged by modifying the objective function (as shown in eq. 2), this paper does not focus on classifying attacks by their goals. Instead, we adopt a more practical approach, categorizing prior attacks based on their motivations (see section 5), including improvements in transferability (section 5.1), physical robustness (section 5.2), stealthiness (section 5.3), and generation speed (section 5.4) of AEs.
@@ 3.3. Victim Models
In defense strategies, image transformation and denoising techniques are commonly used to detect and neutralize adversarial perturbations. In image transformation, methods such as bit-depth compression (Xu, 2017; Guo et al., 2017; Liu et al., 2019), smoothing (Xu, 2017), cropping (Guo et al., 2017), scaling (Guo et al., 2017), and padding (Xie et al., 2017a) are frequently employed. Additionally, JPEG (Guo et al., 2017) adds further steps of image reassembling and quilting. For denoising, (Naseer et al., 2020; Liao et al., 2018; Jia et al., 2019) trained U-Net autoencoders to purify adversarial perturbations, while DiffPure (Nie et al., 2022) utilizes a diffusion model to filter out adversarial noise.
@@ 3.5. Metrics
As shown in Fig. 2, we summarize the metrics commonly used in adversarial attack evaluations. For assessing attack effectiveness, the Attack Success Rate (ASR) and confidence scores from models are the most frequently used. The ASR is defined as the ratio of successful AEs to total test samples, while the average confidence score reflects the mean model confidence on the successful AEs. To evaluate transferability, studies like (Kurakin et al., 2016; Liu et al., 2016; Papernot et al., 2016a, 2017) use the Transfer Rate, which is the ratio of effective transferred AEs to the total test set. For assessing physical robustness, (Kurakin et al., 2018) introduced the destruction rate, defined as the proportion of AEs that lose their adversariality after undergoing physical transformations such as printing or photographing.
For evaluating invisibility, common metrics include the size of perturbations applied to samples and the range of distortions, such as the number of pixels that can be disturbed. Additionally, (Papernot et al., 2016b) introduced hardness metrics and adversarial distances to assess the ease of generating effective AEs. The hardness metric is defined as the integral of the curve formed by the average distortion against the success rate at a given success level. Conversely, the adversarial distance represents the complement of the proportion of pixels that positively contribute to misclassifying the original sample as the target class. In both metrics, a smaller value indicates that it is easier to generate AEs.
For evaluating model robustness, (Moosavi-Dezfooli et al., 2016) proposed Padv, which measures the average distance from test samples to the nearest decision boundary. (Tramèr et al., 2017b, a) argue that the size of the adversarial subspace impacts robustness, proposing the use of the number of dimensions in this space as a metric (with a greater number indicating poorer robustness). A larger adversarial subspace makes it easier to shift original samples into the adversarial domain, leading to decreased model robustness. Additionally, (Fawzi et al., 2016) employ the perturbation norm as another metric for assessing model robustness.

### v2

@@
With the advent of Large Vision-Language Models (LVLMs), new attack vectors, such as cognitive bias, prompt injection, and jailbreaking, have emerged. Understanding these attacks promotes system robustness improvement and neural networks demystification. However, existing surveys often target attack taxonomy and lack in-depth analysis like 1) unified insights into adversariality, transferability, and generalization; 2) detailed evaluations framework; 3) motivation-driven attack categorizations; and 4) an integrated perspective on both traditional and LVLM attacks. This article addresses these gaps by offering a thorough summary of traditional and LVLM adversarial attacks, emphasizing their connections and distinctions, and providing actionable insights for future research.
@@ 1. Introduction
In 2014, Szegedy et al. discovered that small changes to inputs could cause deep neural networks (DNNs) to make drastically different predictions (Szegedy et al., 2014). These carefully crafted inputs were termed adversarial examples (AEs), revealing the vulnerability of deep learning (DL) systems and marking the beginning of adversarial attack and defense research.
1.1. Background
Major Issues of AI Security.
Much like cybersecurity, the security of DL can be classified into three dimensions: availability, confidentiality, and integrity (see Fig. 1). Attacks on availability aim to exhaust system resources (e.g., DDoS, Flood, or energy-latency attacks targeting GPU usage (Gao et al., 2024b)). Attacks on confidentiality involve stealing internal data, such as model parameters or training sets, via model or data extraction (Tramèr et al., 2016). Depending on whether the target is individual data or statistical information, data extraction can be further divided into membership inference (Shokri et al., 2017) or model inversion (Fredrikson et al., 2015). Integrity is primarily threatened by adversarial attacks. Broadly, adversarial attacks fall into poisoning and evasion types. Poisoning attacks target the training sets by injecting malicious data (data poisoning) or the model itself by altering structures or parameters (model poisoning), corrupting the prediction behavior. Evasion attacks occur during inference, where subtle input perturbations mislead the model into incorrect outputs. A subset of poisoning, backdoor attacks, implant hidden triggers into the model, making it misclassify only inputs with specific patterns, thus enhancing stealthiness.
Despite differing goals, these attack types are interrelated. For example, AE generation methods can create poisoned data (Shafahi et al., 2018; Huang et al., 2020). Adversarial training can mitigate backdoor effects (Gao et al., 2023b; Wei et al., 2023), though some researchers argue it lowers the barrier for implanting backdoors (Weng et al., 2020). Recently, Yin et al. found that backdoored models are harder to convert to AEs using triggered inputs, making trigger detection possible by monitoring optimization iterations (Yin et al., 2025). AEs expose the cognitive gap between humans and DL models, manifesting as various attacks depending on specific issues.
This paper focuses on evasion attacks, referring to them as adversarial attacks for simplicity. While early studies also targeted models on traditional learning theories like SVMs, decision trees, and kNNs (Biggio et al., 2013; Papernot et al., 2017, 2016b), most current research focuses on DNNs. Over the past decade, adversarial attacks have evolved from white-box to black-box settings, from prioritizing attack success to balancing stealthiness, robustness, efficiency, and interpretability, and from single-modality classification to multimodal tasks. The remainder of this paper provides a detailed review of this progress.
1.2. Contribution
@@
[Discussions (section 2.1.2), draw=harvestgold, fill=harvestgold!15, text width=3.2cm, text=black]
@@
[Efficiency (section 5.4), draw=majorblue, fill=majorblue!15, text width=3cm, text=black]
Applications and Modalities (section 6), draw=brightlavender, fill=brightlavender!15, text width=3.5cm, text=black
@@
Adversarial attacks manipulate inputs to compromise model integrity, posing significant security threats during machine learning inference. These attacks affect critical applications such as facial recognition (Sharif et al., 2016; Komkov and Petiushko, 2021; Zheng et al., 2023), pedestrian detection (Thys et al., 2019), autonomous driving (Eykholt et al., 2018; Chen et al., 2019; Song et al., 2018), and automated checkout systems (Liu et al., 2020), with severe implications for system security. To improve robustness and safeguard these applications, researchers have pursued extensive investigations, as demonstrated by competitions like NIPS 2017 (NIPS, 2017) and GeekPwn CAAD 2018 (GeekPwn, 2018). A comprehensive understanding of the evolution of adversarial attacks is essential for developing more effective defenses, especially in the Large Vision-Language Models (LVLMs) context. However, classical surveys often fail to capture the latest advancements (Yuan et al., 2019; Akhtar and Mian, 2018; Akhtar et al., 2021), while recent surveys tend to focus on specific areas (Wei et al., 2024b; Badjie et al., 2024; Wang et al., 2024b; Jin et al., 2024; Liu et al., 2024a) or lack thorough summaries (Costa et al., 2024). This paper differentiates itself from existing surveys in several key aspects:
Key Concepts Extraction (section 2). Adversariality, transferability, and generalization are critical traits of AEs that inform design objectives and motivations. This paper fills gaps in previous works by summarizing the causes of adversariality and transferability (section 2.1.1 and section 2.2.1), the roles of AEs (section 6), the properties of transferability (section 2.2.2), and the different types of generalization (section 2.3), which are often overlooked in existing literature.
Motivation Emphasis in Classification (section 4 and section 5). Motivation drives the achievement of goals, which often vary depending on the attacker’s knowledge level and context. As illustrated in Fig. 4, we first categorize attack methods into two attributes based on different dimensions, then further classify them under the second attribute according to varying levels of knowledge, summarizing their design motivations.. Unlike previous works that primarily classify attacks by knowledge levels, we provide a deeper analysis of the motivations behind them.
@@
Given that adversarial attacks have entered the era of LVLMs, this paper categorizes attack methods into traditional and LVLM adversarial attacks based on this pivotal shift. LVLMs are characterized by large training datasets, model capacity, multiple security defenses, and diverse input modalities, in stark contrast to traditional adversarial attacks. As shown in Fig. 2, since classification-based attacks remain the most representative traditional methods, section 4 and section 5 focus on this category. section 7 introduces LVLM attacks, followed by the discussion of the applications and modality extensions (section 6). This paper provides a comprehensive overview of adversarial attack developments, with key contributions:
@@
Identifying emerging attack paradigms and potential vulnerabilities in LVLMs (section 7.1).
@@ 2.1.1. Why Do Adversarial Examples Exist?
Understanding the underlying reasons for AEs is essential for designing stronger attacks and more robust systems. Here, we summarize the hypotheses outlined in previous literature:
Linear Nature of Neural Networks (Goodfellow et al., 2014; Carlini and Wagner, 2017; Papernot et al., 2016b; Tramèr et al., 2017b; Gao et al., 2020b, a). Despite nonlinear activations, DNNs exhibit linear behavior in high-dimensional spaces, contributing significantly to the existence of AEs. Consider a linear model : perturbing the input with can drastically alter predictions when aligns with the direction of the model parameters (Goodfellow et al., 2014). This insight laid the foundation for FGSM (Goodfellow et al., 2014), a seminal adversarial attack. Similarly, CW attack (Carlini and Wagner, 2017) empirically supported the linearity hypothesis by demonstrating that interpolated samples between original inputs and AEs correlate strongly with network logits.
Blind Spots in High-Dimensional Space or Model Overfitting. Due to the limitations of training datasets that can’t fully cover the entire input domain (Papernot et al., 2016b), blind spots may arise (Sharif et al., 2016; Szegedy et al., 2014) or lead to overfitting (Nguyen et al., 2015). (Tanay and Griffin, 2016; Shamir et al., 2021) further proposed the manifold hypothesis to explain such overfitting: When fitting model to data’s low-dimensional manifold (data subspace), the decision boundary develops dimpled local structures (Shamir et al., 2021) or tilts toward low-variance directions of the data (Tanay and Griffin, 2016). This overproximity to data forces samples to be easily separable from the boundary in directions perpendicular to the manifold, thereby generating AEs.
Large Gradient Around Decision Boundaries (Papernot et al., 2016c). This indicates that small perturbations of data points can lead to significant changes in predictions, with points near decision boundaries being potential AEs. This concept also supports the motivations behind the CWA (Chen et al., 2023d) and the RAP (Qin et al., 2022), which enhance transferability by encouraging AEs to converge toward flatter regions—flatter regions contribute to better generalization (Hao et al., 2019).
@@
Additionally, (Szegedy et al., 2014; Sharif et al., 2016; Tabacof and Valle, 2016) suggest that AEs manifest as low-probability, high-density “pockets” within high-dimensional manifolds, rendering them difficult to acquire through random sampling.
2.1.2. Discussions of the Hypotheses
The cause of AEs remains an open question, with a few refutations of hypotheses. For instance, some researchers contend that the linear hypothesis may not hold when perturbations are substantial (Liu et al., 2016; Dong et al., 2018) or when facing specially trained linear classifiers (Tanay and Griffin, 2016). Regarding the “pockets” hypothesis, Tanay et al. (Tanay and Griffin, 2016) argue that Szegedy et al.’s analogy with rational numbers fails to explain the formation of “pockets.” About the high-frequency hypothesis, some studies reveal that networks are also vulnerable to low-frequency information (Guo et al., 2018; Sharma et al., 2019). Notably, although the linear assumption is questioned, its derived attack, FGSM, is widely used and improved, demonstrating its practical validity.
@@ 2.2.2. Characteristics of Transferability
Through analysis and experiments, researchers have gained valuable insights into transferability. For example, (Kurakin et al., 2018, 2016) argue that the transferability of iterative methods is inferior to that of single-step methods, while (Zhou et al., 2018) contend that properly guided gradients can enable iterative methods to achieve good transferability. Additionally, (Kurakin et al., 2016) found that transferability may be inversely proportional to adversariality. By comparing model-aggregated samples (Kurakin et al., 2016), it was shown that universal samples can further enhance transferability (Moosavi-Dezfooli et al., 2017). Moreover, (Kurakin et al., 2016) discovered that adversarial training111Adversarial training is a technique that enhances the robustness of models by incorporating AEs into the training process. with highly transferable samples improves model robustness, whereas (Tramèr et al., 2017b) suggested that a higher dimensionality of adversarial subspaces leads to a greater intersection between the adversarial subspaces of two models, resulting in improved transferability.
@@ 2.3. Generalization
Cross-Environment (Physical Robustness). This generalization allows AEs to maintain their adversariality across different environments (Kurakin et al., 2018; Eykholt et al., 2018; Thys et al., 2019), including various distances, angles, lighting, and device constraints, etc. This phenomenon is often referred to as physical robustness.
@@
Cross-Corpus. This imparts general adversarial semantics to images by alignment across malicious corpus.
Both Cross-Prompt (Bailey et al., 2023; Lu et al., 2024; Luo et al., 2024a) and Cross-Corpus (Qi et al., 2024; Wang et al., 2024e; Ying et al., 2024a) can achieve transfer effects across prompts, but their implementation methods differ. Cross-Corpus aligns the outputs of the LVLM with malicious corpora, enabling perturbations to develop general adversarial semantics and ultimately achieve cross-prompt adversariality. In contrast, Cross-Prompt aggregates perturbations directly across a set of prompts to facilitate generalization.
@@ 3. Problem Setting
In this section, we introduce the threat model and define adversarial attacks in section 3.1 and section 3.2, respectively. This is followed by a discussion of relevant evaluation frameworks in section 3.3, section 3.4, and section 3.5. Since some existing victim models are protected by defense strategies, section 3.3 will also address these defense strategies against adversarial attacks.
@@ 3.1.2. Attacker Goals
Traditional taxonomies divide attack goals into targeted and untargeted attacks. Targeted attacks aim to force the model to produce predictions specified by the attacker, while untargeted attacks require only that the model’s predictions deviate from the correct output. Since targeted and untargeted attacks can be easily interchanged by modifying the objective function (as shown in eq. 2), this paper does not focus on classifying attacks by these goals. Instead, we adopt a more practical way, typing prior attacks based on motivations, including elevation in transferability (section 5.1), physical robustness (section 5.2), stealthiness (section 5.3), generation efficiency (section 5.4), and interpretability (section 5.5).
3.2. Problem Definition
Here, represents the image input modified by , and denotes the desired model output from the attacker. The goal is to identify an attack function for effective input modification. Additionally, data types such as videos and point clouds can also serve as visual inputs. The form of varies depending on the task: for classification, it yields labels; for detection, it produces bounding boxes; and for multimodal tasks, it generate textual responses. Based on the attack intention, targets can be categorized into targeted and untargeted attacks. In targeted attacks, is constrained to be as close as possible to the attacker-specified target output ; in untargeted attacks, must be as far as possible from the original model output . The optimization paradigm for the attack can be expressed as follows:
Here, denotes the distance function. Adversariality is formed through eq. 2, and the stealthiness is discussed in section 5.3.
@@ 3.3. Victim Models
In defenses, image transformation and denoising techniques are commonly used to detect and neutralize perturbations. In image transformation, methods such as bit-depth compression (Xu, 2017; Guo et al., 2017; Liu et al., 2019), smoothing (Xu, 2017), cropping (Guo et al., 2017), scaling (Guo et al., 2017), padding (Xie et al., 2017a), and reassembling (Guo et al., 2017) are frequently employed. For denoising, (Naseer et al., 2020; Liao et al., 2018; Jia et al., 2019) trained U-Net autoencoders to purify adversarial perturbations, while DiffPure (Nie et al., 2022) utilizes a diffusion model. Besides, selective classifiers help filter suspicious samples, enhancing robustness through prediction rejection (Goldwasser et al., 2020) or retraining (Blum et al., 2022).
@@ 3.5. Metrics
As shown in Fig. 3, we summarize key metrics for adversarial attack evaluations. Attack effectiveness is often measured via Attack Success Rate (ASR) or confidence scores. ASR is the percentage of successful AEs, while confidence scores reflects model confidence on them. Transfer rate quantifies transferability, representing the proportion of AEs effective on different models (Kurakin et al., 2016; Liu et al., 2016; Papernot et al., 2016b, 2017). For physical robustness, the destruction rate measures the proportion of AEs losing adversariality after physical transformations like printing or photographing (Kurakin et al., 2018).
Stealthiness is commonly evaluated by perturbation size ( norms222 normalization, or box constraints (Carlini and Wagner, 2017), typically implemented by (where is the budget). norms serve both as evaluation metrics to gauge AE stealthiness and as optimization objectives during generation (section 5.3). When used for optimization, norms 1) constrain perturbed pixel values within a valid color range (Carlini and Wagner, 2017) and 2) limit perturbation magnitude based on attack requirements. ) and distortion range (e.g., the number of affected pixels). Among norms, the norm, which counts altered pixels, has higher computational complexity, while and norms are more sensitive to outliers. These norms are often used in tandem but fail to reflect human perception, making them unreliable for assessing perturbation visibility (Jordan et al., 2019; Wang et al., 2021a; Hameed and Gyorgy, 2021; Karli et al., 2022; Clements and Lao, 2022). Consequently, perception-based metrics like SSIM and LPIPS have been introduced. Additionally, (Papernot et al., 2016c) proposed hardness metrics and adversarial distances to gauge the difficulty of generating AEs. The hardness metric calculates the integral of the curve plotting average distortion against the success rate at a specified threshold, while adversarial distance measures the proportion of pixels not contributing to misclassification. In both cases, lower values indicate easier AE generation.
To assess model robustness, (Moosavi-Dezfooli et al., 2016) propose Padv, calculating the average distance from test samples to the nearest decision boundary. (Tramèr et al., 2017b, a) argue that adversarial subspace dimensionality inversely correlates with robustness—higher dimensions imply greater vulnerability, as expanded subspaces ease the generation of AEs. Additionally, (Fawzi et al., 2016) employ the perturbation norm as another metric for assessing model robustness.